# 1]Equal contributions

### v1

@@ 6.2 Power-Law Scaling Across Data Modalities
The primary results of the experiments are presented in Figure 2 depicting perplexity as a function of training tokens for seven different data modalities. In each plot, both the x-axis (Training Tokens) and y-axis (Perplexity on hold out evaluation set) are on a logarithmic scale. Across all seven conditions—General Domain, Item-Text, CF Both Seen, CF One Unseen, CF Both Unseen, UIH OOD, and UIH Full Graph—the empirical data points exhibit a remarkably consistent trend. The fitted scaling law parameters ($\ell(D)=L_{\infty}+AD^{-\alpha}$ with ideal loss ($L_{\infty}$) and the scaling exponent ($\alpha$)) are summarized in Table 4, with training loss in Figure 6(a).
@@
This quantitative summary reveals a deeper structure in the model’s learning process. The scaling exponent, $\alpha$, can be interpreted as a measure of the "pedagogical efficiency" of each data modality. In the log-log plot, $\alpha$ represents the slope of the line; a larger $\alpha$ value corresponds to a steeper downward slope, signifying that the model’s perplexity decreases more rapidly for each unit of training data consumed. The results in Table 1 show a clear hierarchy of pedagogical efficiency. The Synthetic User Interaction History (UIH) data types exhibit the highest scaling exponents ($\alpha\approx 0.55$), rendering them the most efficient for training. This is followed by the Collaborative Filtering (CF) data types ($\alpha\approx 0.35$), which are themselves significantly more efficient than the Item-Text alignment data ($\alpha\approx 0.15$). The General Domain data, serving as a baseline, is the least efficient ($\alpha\approx 0.025$). This could be due to the pre-trained checkpoint, which is used as initial state of the continual pretraining (CPT), already contains knowledge of general domain, CPT on general domain won’t add new general domain knowledge effectively.
Model with 0.6B parameters has largest $\alpha$ over all sizes as the initial checkpoint has least amount of general domain knowledge. Larger model generally shows larger $\alpha$ for CF and UIH domain, showing their better capabilities in learning the new knowledge; however their $L_{\infty}$ are similar, which suggests even 0.6B parameters sufficient to for CF and UIH domain of this dataset.

### v2

@@ 6.2 Power-Law Scaling Across Data Modalities
The primary results are presented in Figure 2, which depicts perplexity as a function of training tokens for seven data modalities. Both axes use logarithmic scales. Across all conditions—General Domain, Item-Text, CF Both Seen, CF One Unseen, CF Both Unseen, UIH OOD, and UIH Full Graph—the empirical data points follow remarkably consistent power-law trends.
6.2.1 Per-Model Data Scaling Analysis
We first fit per-model scaling laws of the form $\ell(D)=L_{\infty}+AD^{-\alpha}$, where $L_{\infty}$ is the ideal loss and $\alpha$ is the scaling exponent measuring the rate of perplexity reduction per unit of training data. The results are summarized in Table 5.
The scaling exponent $\alpha$ reveals a clear hierarchy of data efficiency across modalities. UIH exhibits the strongest scaling ($\alpha\approx 0.45$$0.59$), followed by CF ($\alpha\approx 0.28$$0.36$), Item-Text ($\alpha\approx 0.13$$0.21$), and General Domain ($\alpha\approx 0.02$$0.03$). The near-saturation of General Domain is expected: the pretrained Qwen3 checkpoint already encodes substantial general-domain knowledge, leaving little room for improvement through continued pretraining on similar data.
Two model-size-dependent patterns emerge. First, larger models exhibit higher $\alpha$ for CF and UIH, indicating that increased model capacity improves the ability to absorb new recommendation knowledge from data. Second, despite this trend, $L_{\infty}$ remains nearly constant across model sizes for CF and UIH, suggesting that even 0.6B parameters are sufficient to represent the underlying patterns in these domains at convergence—the bottleneck is data coverage, not model capacity.
6.2.2 Joint Scaling Law over Model Size and Data
To disentangle the contributions of model size ($N$) and training data ($D$), we fit joint scaling laws of the form $\ell=E+A\cdot N^{-\alpha}+B\cdot D^{-\beta}$, where $E$ is the irreducible loss, $\alpha$ governs model-size scaling, and $\beta$ governs data scaling:
$\ell_{\text{general}}$ $=0.79+16500\,N^{-0.511}+3.85\,D^{-0.048}$ (2)
$\ell_{\text{item-text}}$ $=0.473+5720\,N^{-0.511}+4.04\,D^{-0.070}$ (3)
$\ell_{\text{cf}}$ $=0.193+21\,N^{-0.277}+6.79\,D^{-0.148}$ (4)
$\ell_{\text{UIH}}$ $=0.514+1.89\,N^{-0.138}+63.9\,D^{-0.272}$ (5)
The joint fit yields several findings that complement and deepen the per-model analysis.
Text and recommendation domains exhibit opposite scaling bottlenecks.
There is a striking inverse relationship between $\alpha$ (model-size exponent) and $\beta$ (data exponent) across domains:
  • General/Item-Text: high $\alpha$ (0.511), low $\beta$ (0.048–0.070). These domains are model-size dominant—performance improves primarily by scaling up model capacity.
  • UIH: low $\alpha$ (0.138), high $\beta$ (0.272). This domain is data dominant—performance improves primarily by increasing data coverage.
  • CF: intermediate on both axes ($\alpha=0.277$, $\beta=0.148$), representing a balanced regime.
This pattern is consistent with the per-model observation that $L_{\infty}$ for CF and UIH varies little across model sizes: the joint fit explains this through the small $A$ coefficients for these domains (CF: 21, UIH: 1.89), compared to text domains (General: 16,500, Item-Text: 5,720). The model-size-dependent loss term $AN^{-\alpha}$ contributes minimally for recommendation domains, confirming that their bottleneck lies in data rather than model capacity.
For text domains, the high $\alpha$ but low $\beta$ reflects two complementary factors: (1) the pretrained checkpoint already contains extensive general-domain knowledge, making additional text data marginally redundant; and (2) language modeling involves complex compositional patterns that inherently benefit from increased model capacity.
Compute-optimal allocation differs drastically by domain.
Given a fixed compute budget $C=6ND$, minimizing $\ell$ subject to this constraint yields the optimality condition $\alpha A\cdot N^{-\alpha}=\beta B\cdot D^{-\beta}$, requiring that the marginal loss reduction per FLOP be equal for model size and data. The coefficients $\alpha A$ and $\beta B$ (Table 4) reveal sharply different optimal strategies: General and Item-Text domains should allocate compute overwhelmingly toward model size, while UIH should prioritize data collection. Concretely, halving the data-dependent loss term for General requires $2^{1/0.048}\approx 1.8\times 10^{6}$ times more data, whereas for UIH it requires only $2^{1/0.272}\approx 13$ times more data.
Table 4: Compute-optimal trade-off coefficients. Minimizing loss $\ell=E+AN^{-\alpha}+BD^{-\beta}$ subject to compute budget $C=6ND$ yields the optimality condition $\alpha A\cdot N^{-\alpha}=\beta B\cdot D^{-\beta}$. The ratio of $\alpha A$ to $\beta B$ determines whether compute is best allocated to model size or data.
Domain $\alpha A$ $\beta B$ Implication
General 8,432 0.185 Overwhelmingly model-size dominant
Item-Text 2,923 0.283 Strongly model-size dominant
CF 5.82 1.005 Moderately model-size dominant
UIH 0.261 17.38 Strongly data dominant
Irreducible loss reflects intrinsic predictability.
The irreducible loss $E$ captures the entropy floor of each domain:
  • CF ($E=0.193$): Most predictable, consistent with the low-rank structure of user-item co-occurrence matrices.
  • Item-Text ($E=0.473$): Moderately predictable, reflecting the constrained vocabulary and repetitive structure of item descriptions.
  • UIH ($E=0.514$): User behavior sequences exhibit moderate stochasticity—users are somewhat predictable but not deterministic.
  • General ($E=0.79$): Natural language has the highest inherent entropy, as expected.
CF and UIH capture fundamentally different recommendation signals.
Despite both encoding user behavior, CF and UIH exhibit distinct scaling profiles. CF scales better with model size ($\alpha=0.277$ vs. $0.138$) but less with data ($\beta=0.148$ vs. $0.272$). This distinction reflects their structural differences: CF encodes global user-item co-occurrence patterns, which benefit from model capacity to learn richer latent factors, while UIH encodes sequential, individual-level interaction trajectories, where the primary bottleneck is observational coverage—seeing enough diverse user histories—rather than model expressiveness.
Consistency between the two analyses.
The per-model and joint scaling analyses yield consistent conclusions through complementary lenses. The per-model fits establish the data-efficiency hierarchy (UIH >> CF >> Item-Text >> General), while the joint fit reveals the complementary model-size dimension and explains why the hierarchy exists: domains where data scaling is efficient (high $\beta$) tend to have weak model-size scaling (low $\alpha$), and vice versa. The per-model observation that $L_{\infty}$ is stable across model sizes for CF and UIH is explained by the small $A$ coefficients in the joint fit. However, the per-model analysis also reveals that $\alpha_{\text{data}}$ increases with model size for CF and UIH, suggesting a model-data interaction that the additive joint formulation does not fully capture—an avenue for future investigation.